### v1

@@ Abstract
This paper analyzes and explicitly solves a class of long-term average impulse control problems and a related class of singular control problems. The underlying process is a general one-dimensional diffusion with appropriate boundary behavior. The model is motivated by applications such as the optimal long-term management of renewable resources and financial portfolio management. A large class of admissible policies is identified over which the agent seeks to maximize her long-term average reward, consisting of a running reward and income from either discrete impulses or singular actions. The long-term expected total reward and its relation to overtaking optimality is also considered. Sensitivity analysis with regard to the parameters of the impulse control model are performed. Key connections between the impulse and singular control problems are displayed.
@@ 1 Introduction
$\kappa^{R}:=\limsup_{t\to\infty}t^{-1}\mathbb{E}\Bigg[\sum_{k=1}^{\infty}I_{\{\tau_{k}\leq t\}}Y_{k}\Bigg].$ (1.3)
@@
$J(R;p,K,\gamma):=J(R):=\liminf_{t\to\infty}t^{-1}\mathbb{E}\bigg[\int_{0}^{t}\gamma\,c(X^{R}(s))\,ds+\sum_{k=1}^{\infty}I_{\{\tau_{k}\leq t\}}(pY_{k}-K)\bigg].$ (1.4)
In the context of harvesting problems, the function $\gamma\,c$ can represent the utility derived from maintaining desirable state values $X^{R}(s)$ at time $s$, as well as the state’s contribution to the overall ecosystem’s stability. For example, the function $\gamma\,c$ can be used to model running carbon credits for managing large tracts of forest.
The positive fixed cost for an intervention in (1.4) implies that the optimal policy involves discrete interventions rather than continuous adjustments, ensuring effective product management while maximizing the overall profit rate.
@@
$\widehat{J}(Z;p,\gamma):=\widehat{J}(Z):=\liminf_{t\to\infty}t^{-1}\mathbb{E}_{x_{0}}\bigg[\int_{0}^{t}\gamma\,c(X^{Z}(s))\,ds+p\,Z(t)\bigg].$ (1.6)
@@
The long-term average optimal impulse control problem is addressed in Section 4. Conditions are given for the existence and uniqueness of an optimal $(w,y)$-policy. Furthermore, it is shown that the “impulse reward potential” function $G$ (defined up to a constant in (4.10)) along with the maximal value $F^{*}$ solve the associated system of quasi-variational inequalities (4.12). This, in turn, establishes the optimality of the corresponding $(w,y)$-policy. A brief digression interprets a specific shift of $G$ as the value function of a related infinite-horizon impulse control problem and relates this to the concept of overtaking optimality.
@@ 2 Formulation
$s(x):=\exp\bigg\{-\int_{x_{0}}^{x}\frac{2\mu(y)}{\sigma^{2}(y)}dy\bigg\},\qquad m(x)=\frac{2}{\sigma^{2}(x)s(x)},\qquad x\in(a,b),$ (2.1)
where $x_{0}\in{\mathcal{I}}$ is as in (1.1) and is an arbitrary point, which will be held fixed.
@@
and the potential function $\xi$ defined by
@@
The next proposition highlights implications of Condition 2.1 and introduces the useful identity (2.5), which will be crucial for later calculations. The proof of Proposition 2.2 is placed in Appendix A.1 to preserve the flow of the presentation.
@@
where $\xi^{-}$ denotes the negative part of the function $\xi$, and for each $n\in\mathbb{N}$, $\beta_{n}:=\inf\{t\geq 0:X^{R}(t)\notin(a_{n},b_{n})\}$, in which $\{a_{n}\}\subset{\mathcal{I}}$ is a decreasing sequence with $a_{n}\to a$ and $\{b_{n}\}\subset{\mathcal{I}}$ is an increasing sequence with $b_{n}\to b$.
@@
We finish this section with some elementary but important observations concerning the potential function $\xi$ defined in (2.3) and its companion
$g(x)$ $:=\int_{x_{0}}^{x}\int_{a}^{v}c(u)\,dM(u)\,dS(v),\qquad x\in{\mathcal{I}},$ (2.9)
@@
$\mathbb{E}_{w}[\tau_{y}]$ $=\int_{w}^{y}S[u,y]\,dM(u)+S[w,y]M[a,w]=B\xi(w,y)=\xi(y)-\xi(w),$ (2.13)
$\mathbb{E}_{w}\bigg[\int_{0}^{\tau_{y}}c(X_{0}$ $(s))ds\bigg]=\int_{w}^{y}c(u)S[u,y]\,dM(u)+S[w,y]\int_{a}^{w}c(u)\,dM(u)=Bg(w,y),$ (2.14)
@@
$\mathbb{E}_{w}\bigg[\int_{0}^{\tau_{y}}\mu(X_{0}(s))ds\bigg]=\int_{w}^{y}\mu(u)S[u,y]\,dM(u)+S[w,y]\int_{a}^{w}\mu(u)\,dM(u)=B\mathfrak{i}(w,y).$ (2.15)
@@ 3 $(w,y)$-Impulse Policies
$\kappa^{R^{(w,y)}}=\lim_{t\to\infty}t^{-1}\mathbb{E}\Bigg[\sum_{k=1}^{\infty}I_{\{\tau_{k}\leq t\}}(X^{(w,y)}(\tau_{k}-)-X^{(w,y)}(\tau_{k}))\Bigg]=\frac{y-w}{\xi(y)-\xi(w)}=:{\mathfrak{z}}(w,y).$ (3.2)
@@
$y-w$ $=\int_{w}^{y}\bigg(\int_{a}^{x}\mu(u)m(u)du\bigg)s(x)dx=\varrho^{-1}\int_{a}^{b}\mu(u)\nu(u)du.$
@@
$\mathbb{E}[\xi(X(t\wedge\beta_{n}))]$ $=\xi(x_{0})+\mathbb{E}\Bigg[\int_{0}^{t\wedge\beta_{n}}A\xi(X(s))ds+\sum_{k=1}^{\infty}I_{\{\tau_{k}\leq t\wedge\beta_{n}\}}[\xi(X(\tau_{k}))-\xi(X(\tau_{k}-))]\Bigg]$
$=\xi(x_{0})+\mathbb{E}[t\wedge\beta_{n}]+(\xi(w)-\xi(y))\mathbb{E}\Bigg[\sum_{k=1}^{\infty}I_{\{\tau_{k}\leq t\wedge\beta_{n}\}}\Bigg]$
@@
$\lim_{t\to\infty}\lim_{n\to\infty}t^{-1}\mathbb{E}[\xi(X(t\wedge\beta_{n}))]$ $=1+(\xi(w)-\xi(y))\lim_{t\to\infty}t^{-1}\mathbb{E}\Bigg[\sum_{k=1}^{\infty}I_{\{\tau_{k}\leq t\}}\Bigg]$
@@ 4.1 Existence of an Optimal Policy
$=\lim_{y\to b}\bigg(\frac{p(y-w)}{\xi(y)-\xi(w)}-\frac{K}{\xi(y)-\xi(w)}+\frac{\gamma(g(y)-g(w))}{\xi(y)-\xi(w)}\bigg)=\gamma\bar{c}(b).$ (4.1)
@@
Condition 4.1 indicates that there exists at least one $(w,y)$-policy that outperforms the do-nothing policy $\mathfrak{R}$. We now turn to the maximization of $F$ over pairs $(w,y)\in\mathcal{R}$, a nicely structured fractional optimization problem.
@@
With Proposition 4.2 in our hands, we can now consider the impulse reward potential
@@
where $F^{*}=F(w^{*},y^{*})$ is the maximum value of the function $F$ defined in (3.3), and $(w^{*},y^{*})$ is an optimizing pair for $F$.
@@
$G^{\prime}(x)=\xi^{\prime}(x)\bigg(F^{*}-\frac{\gamma g^{\prime}(x)}{\xi^{\prime}(x)}\bigg)=\xi^{\prime}(x)(F^{*}-h_{0}(x)),$
@@
$\lim_{x\to b}\bigg(F^{*}-\frac{\gamma g(x)}{\xi(x)}\bigg)>0.$
@@
$[F^{*}\xi(x)-\gamma g(x)]I_{\{x\geq x_{0}\}}$ $=[F^{*}\xi(x)-\gamma g(x)]I_{\{x\in[x_{0},y_{0}]\}}+\bigg[F^{*}-\frac{\gamma g(x)}{\xi(x)}\bigg]\xi(x)I_{\{x>y_{0}\}}\geq-C.$
@@
$=G(x_{0})+\mathbb{E}_{x_{0}}\bigg[\int_{0}^{t\wedge\beta_{n}}AG(X(s))\,ds+\sum_{k=1}^{\infty}I_{\{\tau_{k}\leq t\wedge\beta_{n}\}}BG(X(\tau_{k}-),X(\tau_{k}))\bigg]$
$=G(x_{0})+F^{*}\mathbb{E}_{x_{0}}[t\wedge\beta_{n}]-\mathbb{E}_{x_{0}}\bigg[\int_{0}^{t\wedge\beta_{n}}\gamma c(X(s))ds+\sum_{k=1}^{\infty}I_{\{\tau_{k}\leq t\wedge\beta_{n}\}}BG(X(\tau_{k}-),X(\tau_{k}))\bigg].$
@@
$\ \qquad=\mathbb{E}_{x_{0}}\bigg[\int_{0}^{t\wedge\beta_{n}}\gamma c(X(s))ds+\sum_{k=1}^{\infty}I_{\{\tau_{k}\leq t\wedge\beta_{n}\}}[G(X(\tau_{k}-))-G(X(\tau_{k}))]\bigg].$
@@
$=\mathbb{E}_{x_{0}}\bigg[\int_{0}^{t}\gamma c(X(s))ds+\sum_{k=1}^{\infty}I_{\{\tau_{k}\leq t\}}[G(X(\tau_{k}-))-G(X(\tau_{k}))]\bigg]$
$\geq\mathbb{E}_{x_{0}}\bigg[\int_{0}^{t}\gamma c(X(s))ds+\sum_{k=1}^{\infty}I_{\{\tau_{k}\leq t\}}[p(X(\tau_{k}-)-X(\tau_{k}))-K]\bigg],$ (4.16)
where the last inequality follows from the second line of (4.11). Also note that the last inequality holds as an equality when the $(w^{*},y^{*})$-policy $R^{(w^{*},y^{*})}$ is used. Now divide both sides by $t$ and then let $t\to\infty$ to obtain
@@
$\geq\limsup_{t\to\infty}\frac{1}{t}\mathbb{E}_{x_{0}}\bigg[\int_{0}^{t}\gamma c(X(s))ds+\sum_{k=1}^{\infty}I_{\{\tau_{k}\leq t\}}[p(X(\tau_{k}-)-X(\tau_{k}))-K]\bigg].$
@@
$F^{*}\geq\limsup_{t\to\infty}\frac{1}{t}\mathbb{E}_{x_{0}}\bigg[\int_{0}^{t}\gamma c(X(s))ds+\sum_{k=1}^{\infty}I_{\{\tau_{k}\leq t\}}[p(X(\tau_{k}-)-X(\tau_{k}))-K]\bigg].$ (4.17)
Moreover, equality holds by the third line of (4.11) when the $(w^{*},y^{*})$-policy $R^{(w^{*},y^{*})}$ is used. This establishes (4.15). Note that $R^{(w^{*},y^{*})}\in{\mathcal{A}}$ thanks to Lemma 3.2. Hence $R^{(w^{*},y^{*})}$ is an optimal policy. This completes the proof. ∎
@@
$\lim_{x\to a}G(x)=\lim_{x\to a}\bigg(F^{*}-\frac{\gamma g(x)}{\xi(x)}\bigg)\xi(x)=\lim_{x\to a}(F^{*}-\gamma c(a))\xi(x)=-\infty.$
@@ 4.2 Uniqueness
$=\frac{m(x)}{M[a,x]^{2}}\bigg[\int_{a}^{\widehat{y}}[r(\widehat{y})-r(u)+r(x)-r(\widehat{y})]\,dM(u)+\int_{\widehat{y}}^{x}[r(x)-r(u)]\,dM(u)\bigg]$

### v2

@@ Abstract
This paper analyzes and explicitly solves a class of long-term average impulse control problems and a related class of singular control problems. The underlying process is a general one-dimensional diffusion with appropriate boundary behavior. The model is motivated by applications such as the optimal long-term management of renewable resources and financial portfolio management. A large class of admissible policies is identified over which the agent seeks to maximize her long-term average reward, consisting of a running reward and income from either discrete impulses or singular actions. In addition to the long-term average objective, we will briefly consider the long-term expected total reward functional and its relation to overtaking optimality. Sensitivity analysis with regard to the parameters of the impulse control model are performed. Key connections between the impulse and singular control problems are displayed.
@@ 1 Introduction
$\kappa^{R}:=\limsup_{t\to\infty}t^{-1}\mathbb{E}\left[\sum_{k=1}^{\infty}I_{\{\tau_{k}\leq t\}}Y_{k}\right].$ (1.3)
@@
$J(R;p,K,\gamma):=J(R):=\liminf_{t\to\infty}t^{-1}\mathbb{E}\left[\int_{0}^{t}\gamma\,c(X^{R}(s))\,ds+\sum_{k=1}^{\infty}I_{\{\tau_{k}\leq t\}}(pY_{k}-K)\right].$ (1.4)
In the context of harvesting problems, the function $\gamma\,c$ can represent the utility derived from maintaining desirable state values $X^{R}(s)$ at time $s$, as well as the state’s contribution to the overall ecosystem’s stability. For example, the function $\gamma\,c$ can be used to model running carbon credits for managing large tracts of forest. The positive fixed cost for each intervention in (1.4) implies that the optimal policy involves discrete interventions rather than continuous adjustments, ensuring effective product management while maximizing the overall profit rate.
@@
$\widehat{J}(Z;p,\gamma):=\widehat{J}(Z):=\liminf_{t\to\infty}t^{-1}\mathbb{E}_{x_{0}}\left[\int_{0}^{t}\gamma\,c(X^{Z}(s))\,ds+p\,Z(t)\right].$ (1.6)
@@
The long-term average optimal impulse control problem is addressed in Section 4. Conditions are given for the existence and uniqueness of an optimal $(w,y)$-policy. Furthermore, it is shown that the “impulse reward potential” function $G$ (defined up to a constant in (4.10)) along with the maximal value $F^{*}$ is a solution of the associated system of quasi-variational inequalities (4.12). This, in turn, establishes the optimality of the corresponding $(w,y)$-policy. The potential $G$ is then used to define the natural solution $\widetilde{G}$ of the system (4.12). A brief digression examines the growth of the long-term expected total reward. A specific shift of $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\widetilde{G}$ is interpreted as the value function of an infinite-horizon growth rate optimization problem. The section concludes by relating the optimal $(w,y)$-policy to the concept of an optimal overtaking policy.
@@ 2 Formulation
    $s(x):=\exp\left\{-\int_{x_{0}}^{x}\frac{2\mu(y)}{\sigma^{2}(y)}dy\right\},\qquad m(x)=\frac{2}{\sigma^{2}(x)s(x)},\qquad x\in(a,b),$ (2.1)
    where $x_{0}\in{\mathcal{I}}$ is as in (1.1) and is an arbitrary point, which will be held fixed. (Note we follow the convention of using the factor $2$ in the definition of $m$.)
  • (b)
@@
    and the (hitting-time) potential function $\xi$ defined by
@@
The next proposition highlights implications of Condition 2.1 and introduces the useful identity (2.5), which will be crucial for both theoretical and numerical calculations. The proof of Proposition 2.2 is placed in Appendix A.1 to preserve the flow of the presentation.
@@
      where $\xi^{-}$ denotes the negative part of the time potential $\xi$, and for each $n\in\mathbb{N}$, $\beta_{n}:=\inf\{t\geq 0:X^{R}(t)\notin(a_{n},b_{n})\}$, in which $\{a_{n}\}\subset{\mathcal{I}}$ is a decreasing sequence with $a_{n}\to a$ and $\{b_{n}\}\subset{\mathcal{I}}$ is an increasing sequence with $b_{n}\to b$.
@@
We finish this section with some elementary but important observations concerning the time potential $\xi$ defined in (2.3) and its (running reward) companion
$g(x)$ $:=\int_{x_{0}}^{x}\int_{a}^{v}c(u)\,dM(u)r\,dS(v),\qquad x\in{\mathcal{I}},$ (2.9)
@@
$\mathbb{E}_{w}$ $[\tau_{y}]=\int_{w}^{y}S[u,y]\,dM(u)+S[w,y]M[a,w]=\xi(y)-\xi(w)=B\xi(w,y),$ (2.13)
$\mathbb{E}_{w}$ $\left[\int_{0}^{\tau_{y}}c(X_{0}(s))ds\right]=\int_{w}^{y}c(u)S[u,y]\,dM(u)+S[w,y]\int_{a}^{w}c(u)\,dM(u)=Bg(w,y),$ (2.14)
@@
$\mathbb{E}_{w}\left[\int_{0}^{\tau_{y}}\mu(X_{0}(s))ds\right]=\int_{w}^{y}\mu(u)S[u,y]\,dM(u)+S[w,y]\int_{a}^{w}\mu(u)\,dM(u)=B\mathfrak{i}(w,y).$ (2.15)
@@ 3 $(w,y)$-Impulse Policies
$\kappa^{R^{(w,y)}}=\lim_{t\to\infty}t^{-1}\mathbb{E}\left[\sum_{k=1}^{\infty}I_{\{\tau_{k}\leq t\}}(X^{(w,y)}(\tau_{k}-)-X^{(w,y)}(\tau_{k}))\right]=\frac{y-w}{\xi(y)-\xi(w)}=:{\mathfrak{z}}(w,y).$ (3.2)
@@
$y-w$ $=\int_{w}^{y}\left(\int_{a}^{x}\mu(u)m(u)du\right)s(x)dx=\varrho^{-1}\int_{a}^{b}\mu(u)\nu(u)du.$
@@
$\mathbb{E}[\xi(X(t\wedge\beta_{n}))]$ $=\xi(x_{0})+\mathbb{E}\left[\int_{0}^{t\wedge\beta_{n}}A\xi(X(s))ds+\sum_{k=1}^{\infty}I_{\{\tau_{k}\leq t\wedge\beta_{n}\}}[\xi(X(\tau_{k}))-\xi(X(\tau_{k}-))]\right]$
$=\xi(x_{0})+\mathbb{E}[t\wedge\beta_{n}]+(\xi(w)-\xi(y))\mathbb{E}\left[\sum_{k=1}^{\infty}I_{\{\tau_{k}\leq t\wedge\beta_{n}\}}\right]$
@@
$\lim_{t\to\infty}\lim_{n\to\infty}t^{-1}\mathbb{E}[\xi(X(t\wedge\beta_{n}))]$ $=1+(\xi(w)-\xi(y))\lim_{t\to\infty}t^{-1}\mathbb{E}\left[\sum_{k=1}^{\infty}I_{\{\tau_{k}\leq t\}}\right]$
@@ 4.1 Existence of an Optimal Policy
$=\lim_{y\to b}\left(\frac{p(y-w)}{\xi(y)-\xi(w)}-\frac{K}{\xi(y)-\xi(w)}+\frac{\gamma(g(y)-g(w))}{\xi(y)-\xi(w)}\right)=\gamma\bar{c}(b).$ (4.1)
@@
Condition 4.1 indicates that there exists at least one $(w,y)$-policy that outperforms the do-nothing policy $\mathfrak{R}$.
We now turn to the maximization of $F$ over pairs $(w,y)\in\mathcal{R}$, a nicely structured fractional optimization problem.
@@
Motivated by Proposition 4.2, we now consider the impulse reward potential
@@
where $F^{*}=F(w^{*},y^{*})$ is the maximum value of the function $F$ defined in (3.3), and $(w^{*},y^{*})$ is an optimizing pair for $F$. The potential $G$ can be used to enlarge the class of admissible policies (see Remark 4.6) and is intimately related to the relative value function for the impulse control problem.
@@
$G^{\prime}(x)=\xi^{\prime}(x)\left(F^{*}-\frac{\gamma g^{\prime}(x)}{\xi^{\prime}(x)}\right)=\xi^{\prime}(x)(F^{*}-h_{0}(x)),$
@@
$\lim_{x\to b}\left(F^{*}-\frac{\gamma g(x)}{\xi(x)}\right)>0.$
@@
$[F^{*}\xi(x)-\gamma g(x)]I_{\{x\geq x_{0}\}}$ $=[F^{*}\xi(x)-\gamma g(x)]I_{\{x\in[x_{0},y_{0}]\}}+\left[F^{*}-\frac{\gamma g(x)}{\xi(x)}\right]\xi(x)I_{\{x>y_{0}\}}\geq-C.$
@@
$=G(x_{0})+\mathbb{E}_{x_{0}}\left[\int_{0}^{t\wedge\beta_{n}}AG(X(s))\,ds+\sum_{k=1}^{\infty}I_{\{\tau_{k}\leq t\wedge\beta_{n}\}}BG(X(\tau_{k}-),X(\tau_{k}))\right]$
$=G(x_{0})+F^{*}\mathbb{E}_{x_{0}}[t\wedge\beta_{n}]-\mathbb{E}_{x_{0}}\left[\int_{0}^{t\wedge\beta_{n}}\gamma c(X(s))ds+\sum_{k=1}^{\infty}I_{\{\tau_{k}\leq t\wedge\beta_{n}\}}BG(X(\tau_{k}-),X(\tau_{k}))\right].$
@@
$\ \qquad=\mathbb{E}_{x_{0}}\left[\int_{0}^{t\wedge\beta_{n}}\gamma c(X(s))ds+\sum_{k=1}^{\infty}I_{\{\tau_{k}\leq t\wedge\beta_{n}\}}[G(X(\tau_{k}-))-G(X(\tau_{k}))]\right].$
@@
$=\mathbb{E}_{x_{0}}\left[\int_{0}^{t}\gamma c(X(s))ds+\sum_{k=1}^{\infty}I_{\{\tau_{k}\leq t\}}[G(X(\tau_{k}-))-G(X(\tau_{k}))]\right]$ (4.16)
$\geq\mathbb{E}_{x_{0}}\left[\int_{0}^{t}\gamma c(X(s))ds+\sum_{k=1}^{\infty}I_{\{\tau_{k}\leq t\}}[p(X(\tau_{k}-)-X(\tau_{k}))-K]\right],$
where the last inequality follows from the second line of (4.11). Now divide both sides by $t$ and then let $t\to\infty$ to obtain
@@
$\geq\limsup_{t\to\infty}\frac{1}{t}\mathbb{E}_{x_{0}}\left[\int_{0}^{t}\gamma c(X(s))ds+\sum_{k=1}^{\infty}I_{\{\tau_{k}\leq t\}}[p(X(\tau_{k}-)-X(\tau_{k}))-K]\right].$
@@
$F^{*}\geq\limsup_{t\to\infty}\frac{1}{t}\mathbb{E}_{x_{0}}\left[\int_{0}^{t}\gamma c(X(s))ds+\sum_{k=1}^{\infty}I_{\{\tau_{k}\leq t\}}[p(X(\tau_{k}-)-X(\tau_{k}))-K]\right].$ (4.17)
This establishes (4.15). On the other hand, we have shown in Section 3 that the $R^{(w^{*},y^{*})}\in{\mathcal{A}}$ with $J(R^{(w^{*},y^{*})})=F^{*}$. Hence $R^{(w^{*},y^{*})}$ is an optimal policy. This completes the proof. ∎
@@
$\lim_{x\to a}G(x)=\lim_{x\to a}\left(F^{*}-\frac{\gamma g(x)}{\xi(x)}\right)\xi(x)=\lim_{x\to a}(F^{*}-\gamma c(a))\xi(x)=-\infty.$
@@ 4.2 Uniqueness
$=\frac{m(x)}{M[a,x]^{2}}\left[\int_{a}^{\widehat{y}}[r(\widehat{y})-r(u)+r(x)-r(\widehat{y})]\,dM(u)+\int_{\widehat{y}}^{x}[r(x)-r(u)]\,dM(u)\right]$